OpenStudy (anonymous):

find the limit as x approaches 0 of (cosx+2x)^1/x

OpenStudy (dumbcow):

first recognize that e^ln(x) = x so \[\large e^{\lim_{x \rightarrow 0} \ln(\cos x +2x)^{1/x}}\] \[\rightarrow \large e^{\lim_{x \rightarrow 0} \frac{\ln(\cos x +2x)}{x}}\] this yields lim 0/0 which is indeterminate use L'Hopitals rule, take derivative of top/bottom \[\rightarrow \large e^{\lim_{x \rightarrow 0} \frac{2-\sin x}{\\cos x +2x}}\] \[= e^2\]

OpenStudy (anonymous):

isn't the derivative of ln(cosx+2x) 1/(cosx+2x) *-sinx +1

OpenStudy (dumbcow):

no derivative of 2x is 2

OpenStudy (anonymous):

oh I forgot to take that derivative properly..

OpenStudy (anonymous):

Thank you so much, I was having trouble finding out what I did wrong.

OpenStudy (dumbcow):

no problem